a.     A proton (q=+5x10-19C) is injected from right to left into a field B of 0.5 T directed toward the upper side of a sheet of paper. If the velocity of the proton is of 4x106m/s, find the magnitude and direction of the magnetic force over the proton.

b.      A wire 50 cm long conducts a current of 1A forming an angle of 35° to the north of a magnetic field directed toward east. Find the magnitude of the magnetic field so that a force of 17N can be exerted over the wire. Find the direction of the force.



1
Expert's answer
2021-11-19T10:46:51-0500

"F= Force\\q= charge\\v= velocity \\B=magnetic\>field\\\theta \> ="

angle between the particle and the magnetic field

Then:


"F=qvBsin\theta"


"F= 5\u00d710^{-19}\u00d74\u00d710^6\u00d70.5 sin90"


"F=1×10^{-12}\>N"


Part 2


"F=\Iota lBsin\theta"


Where

"F= Force\\\Iota=current\\L="length of conductor

"B" =manetic field


"B=\frac{17}{1×0.5×sin35}"


="59.28 T"


Force will be directed upward.